# The Trans-Neptunian Object (119951) 2002 KX14 as revealed by multiple stellar occultations

### v1

@@ Abstract
Context. Trans-Neptunian Objects (TNOs) are icy bodies located in the outer solar system, providing key insights into the primordial conditions of our planetary system. The stellar occultation technique has proven to be an essential tool for studying these distant and faint objects, enabling precise determination of their size, shape, albedo, and improving their orbital parameters. Among these, (119951) 2002 KX14 is a large classical TNO with limited previous observations and unresolved questions regarding its physical properties.
Aims. This study aims to analyze and characterize the TNO (119951) 2002 KX14 through observations of stellar occultations, providing unique measurements of its size, shape and albedo. Measuring shapes is especially relevant as only thirteen TNOs have their projected shapes measured. These results contribute to understanding the physical properties of this object and the broader TNO population.
Methods. Five stellar occultations by 2002 KX14 were observed from 2020 to 2023, involving multiple telescopes across different locations in Europe and the Americas. High-precision astrometry and photometric data were used to predict the occultation paths and extract ingress and egress timings. One of the events was detected from six sites and there are also several close misses, providing precise constraints for an accurate determination of the object’s limb. Then, elliptical fits to the occultations chords allowed for the determination of the object’s shape and area-equivalent diameter. Geometric albedo was calculated by combining the occultation results with published absolute magnitudes.
Results. The five occultations resulted in 15 positive chords, accurately measuring the 2002 KX14’s shape and size. Given that the rotational variability of this body is minimal, we can reasonably assume that the variations are due to albedo features, being the body a Maclaurin spheroid. The projected ellipse has semi-major and semi-minor axes of $241.0\pm 7.2$ km and $157.1\pm 5.2$ km, respectively, corresponding to an average area-equivalent diameter of $389.2\pm 8.7$ km. The geometric albedo was estimated at $11.9\pm 0.7\%$. These values contrast with the $455\pm 27$ km diameter and $9.7^{+1.4}_{-1.3}\%$ albedo reported by Vilenius et al. (2012) and provide a more refined understanding of the object’s physical properties.
@@ 1 Introduction
Beyond Neptune lies a vast and diverse population of Trans-Neptunian Objects (TNOs) [^1] At the time of writing this article, more than 5,000 objects, including the scattered disk and Centaurs, are listed at https://minorplanetcenter.net/iau/lists/TNOs.html and https://minorplanetcenter.net/iau/lists/Centaurs.html., which can be dynamically classified into several groups based on their orbital characteristics (Gladman et al. 2008). On the one hand, the Resonant Objects, such as Plutinos, in a 3:2 resonance with Neptune, follow predictable orbital patterns influenced by the planet’s gravity. On the other hand, the Classical Belt (no resonants), is divided into the “cold population” with low inclinations and nearly circular orbits, and the “hot population” with higher inclinations and eccentricities. Beyond this lies the Scattered Disk, containing objects with highly eccentric and inclined orbits, thought to have been dynamically perturbed by Neptune. Further out, detached objects like Sedna have perihelia that is far from Neptune’s influence, while some may represent the transition to the hypothesized inner Oort cloud.
@@
The remote location of these icy bodies, which shields them from the intense solar radiation that affects objects closer to the Sun, makes them valuable for studying the early physical and dynamical conditions that shaped our planetary system. However, it is challenging given their relatively small size (the largest TNO is Pluto, with a radius of 1188 km (Nimmo et al. 2017)), low brightness, and great heliocentric distances. Moreover, the low surface temperatures of TNOs (30-50 K, e.g. Stern & Trafton (2008)) cause their thermal emission to peak in the far-infrared region of the electromagnetic spectrum, which is blocked by Earth’s atmosphere. This makes space-based observations essential for their study. However, the stellar occultation technique enables ground-based observations and has proven to be one of the most effective methods for studying TNOs to date. This technique has significantly advanced our understanding of these distant icy objects by allowing for highly accurate size measurements, revealing the presence of atmospheres, moons, or rings, and, when combined with independent datasets, determining their geometric albedo (e.g. Hubbard et al. (1988); Sicardy et al. (2011); Ortiz et al. (2017); Assafin et al. (2023)). Moreover, it achieves kilometric precision, making it especially useful for refining their orbital parameters through astrometry and determining a highly accurate shape.
@@
In this work, we focus on the trans-Neptunian object (119951) 2002 KX14, discovered in 2002, which has a semi-major axis of 38.63 au, an eccentricity of 0.04, and an inclination of 0.41°. This object exhibits orbital characteristics—specifically low eccentricity and low inclination—that are typical of the cold classical population. Members of this population are the most dynamically stable objects in the classical belt and are believed to have formed in situ, making them valuable tracers of the early formation history of the trans-Neptunian region (Delsanti & Jewitt 2006). However, the semi-major axis of 2002 KX14 places it outside the classical belt region, generally defined between $\sim$42 and 47 au. Some authors (e.g., Fernández-Valenzuela et al. 2021) have classified it as an inner classical object, a population considered not an extension of the cold classicals but belonging to the low-inclination tail of the hot population (Vilenius et al. 2012). Unlike the cold classicals, the hot population is thought to have been dynamically excited and relocated to its current position during the migration of the giant planets. During this migration, gravitational interactions perturbed their orbits, increasing both inclination and eccentricity. Therefore, studying 2002 KX14 is particularly valuable because its orbital properties place it near the boundary between the cold and hot classical populations, offering insights into the dynamical structure and evolutionary history of the trans-Neptunian region.
@@
Based on thermal measurements with the Herschel Space Telescope, Vilenius et al. (2012) estimated a diameter of $455\pm 27$ km and a geometric albedo ($p_{V})$ of $9.7\begin{subarray}{c}+1.4\\ -1.3\end{subarray}\%$. This diameter is close to the size where hydrostatic equilibrium may no longer be maintained (Tancredi & Favre 2008). Subsequent ground-based observations studied its photometric properties, providing values of its absolute magnitude in the R-band: $H_{R}=4.349\pm 0.124$ (Peixinho et al. 2012), $H_{R}=4.46\pm 0.01$ (Benecchi & Sheppard 2013), and in the V-band: $H_{V}=5.07\pm 0.03$ (Boehnhardt et al. 2014), $H_{V}=4.978\pm 0.017$ (Alvarez-Candal et al. 2016). Also, Alvarez-Candal et al. (2014) recorded a positive occultation, measuring a chord of $415\pm 1$ km and retrieving an equivalent diameter of $365\begin{subarray}{c}+30\\ -21\end{subarray}$ km. Later, Rommel et al. (2020) got two additional chords of $421\pm 3$ km and $431\pm 6$ km. Despite these observations, due to its faint apparent magnitude above 20.5 in the V-band, with rotational variations under 0.05 mag (Benecchi & Sheppard 2013), the shape of 2002 KX14, along with its pole and rotation period is unknown.
@@
In this study, we present the results of five positive occultations, which yield a total of 15 chords on five different dates. These observations have allowed us to determine the object’s shape and geometric albedo. Currently, only thirteen TNOs have projected shapes determined (Ortiz et al. 2012; Benedetti-Rossi et al. 2016; Ortiz et al. 2017; Dias-Oliveira et al. 2017; Benedetti-Rossi et al. 2019; Ortiz et al. 2020; Souami et al. 2020; Santos-Sanz et al. 2022; Vara-Lubiano et al. 2022; Pereira et al. 2023; Rommel et al. 2023; Kretlow et al. 2024), with 2002 KX14 being now the first [^2] Apart from Arrokoth, which was observed up close by the New Horizons mission cold classical object (if that turns out to be the case). The fact that it is close to the critical diameter beyond which hydrostatic equilibrium may no longer be maintained makes determining its shape particularly interesting. Thus, this study contributes to the growing body of knowledge on TNOs, lays the groundwork for future in-depth studies, provides data to enhance the statistics that allow us to contextualize this population more comprehensively, and demonstrates the effectiveness of the occultation technique in characterizing these distant objects.
@@ 2 Observations
Table 1: Occulted stars by 2002 KX14 presented in this work. Stellar right ascension (h m s) and declination (d m s) represent the astrometric coordinates on the International Celestial Reference Frame (ICRF) of the stars, corrected of proper motion, given for the moments of the occultations.
@@
2.1 Stellar occultations
Initial predictions for 2002 KX14 were made using the Numerical Integration of the Motion of an Asteroid (NIMA) solution (Desmars et al. 2015). Subsequently, several observational campaigns were conducted prior to each occultation using the 1.23 m telescope at the Calar Alto Observatory and the 2.0 m Liverpool Telescope located at the Roque de los Muchachos Observatory. These efforts aimed to perform astrometry of this TNO and refine the predicted occultation shadow path, which was communicated to observers for proper deployment in cases for which mobile equipment was used. This approach enabled the observation of five positive occultations in 2020, 2022, and 2023, as described in the following sections.
2.1.1 Occultation A - May 26, 2020
@@
The event was observed from nine different locations, resulting in six positive detections from Borowiec (Poland), Odesa-Mayaki Observatory (Ukraine), Konkoly Observatory (Hungary), Gothard Astrophysical Observatory (Hungary), Millen-Bruch (Germany), Nandrin (Belgium), and two negative detections from Buelach (Switzerland) and Çukurova (Turkey). In addition to this, it was also observed from Tubitak Observatory (Turkey) and Marseilles (France), but the signal-to-noise ratio (S/N) of the data was too low to confirm whether the occultation occurred or not. An additional positive detection was obtained from Strasice (Czechia), but due to technical issues with the data format, it was not possible to determine the ingress and egress times with sufficient accuracy to be useful for this work. Fig.3(a) shows the predicted occultation shadow path along with the position of the observing stations involved in this work. Table 3 expands on the information regarding locations, instrumentation, and the observers involved. Time series observations were obtained and time synchronized using Network Time Protocol (NTP) servers or global positioning systems (GPS) devices. Image acquisition started at least $\sim$5 min before and ended $\sim$5 min after the occultation time. No filters were used to maximize the S/N of the occulted star.
@@
As it is a giant star (see the Hertzsprung–Russell diagram in Fig. 2) $(a,b)=(0.669,0.223)$ according to van Belle (1999). $V$ and $K$ denote the apparent magnitudes of the star in the V-band and K-band, being $V$ = 14.750 mag and $K$ = 10.520 mag (Zacharias et al. 2004). Consequently, the angular diameter is calculated as 0.046 mas, corresponding to 1.26 km at a 2002 KX14’s geocentric distance ($\Delta$) of 38.01 au. As the central wavelength ($\lambda$) is $\sim$600 nm for our CCD/CMOS observations, from Eq. 2, the Fresnel scale of this object is calculated as 1.3 km,
@@
2.1.2 Occultation B - June 23, 2022
On June 23, 2022, an occultation of the star Gaia DR3 4116601049409385472 by 2002 KX14 (Table 1), with a predicted maximum duration of 18.7 seconds, was forecasted to occur at 00:51:33 UT (hereafter ”Occ. B”). The event was observed from four locations in Spain: the Botorrita, Javalambre, La Hita, and Calar Alto observatories. Three of these observatories—Javalambre, La Hita, and Calar Alto—reported positive detections, while Botorrita experienced technical issues (Fig.3(b), Table 3). Time series observations were captured and synchronized using either NTP servers or GPS devices. Image acquisition started approximately $\sim$5 minutes before and concluded $\sim$5 minutes after the predicted event time. No filters were applied to optimize the S/N of the star.
@@ 2.1.3 Occultation C - May 17, 2023
An angular diameter of 0.046 mas of the occulted star was estimated using Eq. 1 (giant star, see Fig. 2), with $V$ = 16.47 mag and $K$ = 10.71 mag from Zacharias et al. (2004). This corresponds to 1.25 km at 2002 KX14’s geocentric distance, $\Delta$, of 37.97 au. The Fresnel scale in this case is 1.3 km. Because the minimum cycle time for these observations is 2.0 seconds, while 2002 KX14’s velocity at the time of occultation was 20.38 km/s, it translates to a distance of 40.76 km. Given that this distance is orders of magnitude greater than that derived from the star’s angular size and Fresnel diffraction effects, stellar angular size and diffraction have a small impact on the derivation of ingress and egress times. Yet, they were accounted for.
2.1.4 Occultation D - May 19, 2023
Two days later, on May 19, 2023, an occultation by 2002 KX14 of the star Gaia DR3 4068793394867887872 (Table 1) with a maximum duration of 21.8 seconds was predicted to happen at 03:08:49 UT (hereafter ”Occ. D”). Observations of this event were conducted from three locations in Brazil, with positive detections reported by the GOA and Sonear3 observatories, while in DogsHeaven the S/N of the data was too low (Fig.3(d), Table 3). Time series data were collected and synchronized using NTP servers. Image capture began approximately $\sim$5 min before the event and continued for about five minutes afterward. To maximize the star’s S/N, no filters were applied.
Using Eq.1 (giant star, see Fig. 2) the occulted star’s diameter was estimated to be 0.032 mas, with magnitudes of $V$ = 16.39 mag and $K$ = 11.36 mag from Zacharias et al. (2004). At 2002 KX14’s geocentric distance of 37.96 au, this angular size corresponds to a physical diameter of 0.89 km. The Fresnel scale is again 1.3 km. Given that the minimum observation cycle time is 6.8 seconds and that 2002 KX14’s velocity during the occultation was 20.83 km/s, this timing translates to a spatial distance of 141.64 km. Thus, the ingress and egress determination uncertainty mainly comes from the photometric accuracy of the occultation light curves. Even so, the star’s angular size and Fresnel diffraction effects were included in the analysis.
2.1.5 Occultation E - July 01, 2023

### v2

# The trans-Neptunian object (119951) 2002 KX14 revealed via multiple stellar occultations

@@ Abstract
Context. Trans-Neptunian objects (TNOs) are icy bodies located in the outer solar system that offer key insights into the primordial conditions of our planetary system. The stellar occultation technique has proven to be an essential tool for studying these distant and faint objects, enabling precise determinations of their size, shape, and albedo, while also improving estimates of their orbital parameters. Among them, (119951) 2002 KX14 is a large classical TNO with limited previous observations and unresolved questions regarding its physical properties.
Aims. This study aims to analyze and characterize the TNO (119951) 2002 KX14 through observations of stellar occultations, providing unique measurements of its size, shape, and albedo. Shape measurements are especially relevant, as only thirteen TNOs have had their projected shapes measured. These results contribute to our understanding of the physical properties of this object and the broader TNO population.
Methods. Five stellar occultations by 2002 KX14 were observed from 2020 to 2023, involving multiple telescopes across different locations in Europe and the Americas. High-precision astrometry and photometric data were used to predict the occultation paths and extract ingress and egress timings. One of the events was detected from six sites and there are also several close misses, providing precise constraints for an accurate determination of the object’s limb. Furthermore, elliptical fits to the occultations chords allowed for the determination of the object’s shape and area-equivalent diameter. The geometric albedo was calculated by combining the occultation results with published absolute magnitudes.
Results. The five occultations resulted in 15 positive chords that allowed us to accurately measure the shape and size of 2002 KX14. Given that the rotational variability of this body is minimal, we can reasonably assume that the variations are due to albedo features, since the body is classed as a Maclaurin spheroid. The projected ellipse has semi-major and semi-minor axes of $241.0\pm 7.2$ km and $157.1\pm 5.2$ km, respectively, corresponding to an average area-equivalent diameter of $389.2\pm 8.7$ km. The geometric albedo was estimated to be $11.9\pm 0.7\%$. These values differ from the $455\pm 27$ km diameter and the $9.7^{+1.4}_{-1.3}\%$ albedo derived from thermal measurements, offering a more refined understanding of the object’s physical properties.
@@ 1 Introduction
Beyond Neptune there lies a vast and diverse population of trans-Neptunian objects (TNOs) [^1] At the time of writing, more than 5,000 objects, including the scattered disk and Centaurs, have been listed at https://minorplanetcenter.net/iau/lists/TNOs.html and https://minorplanetcenter.net/iau/lists/Centaurs.html., which can be dynamically classified into several groups based on their orbital characteristics (Gladman et al. 2008). On the one hand, resonant objects, such as Plutinos, are in a 3:2 resonance with Neptune and follow predictable orbital patterns influenced by the planet’s gravity. On the other hand, the classical belt (nonresonants), is divided into a “cold population,” with low inclinations and nearly circular orbits, and a “hot population,” with higher inclinations and eccentricities. Beyond this lies the scattered disk, containing objects with highly eccentric and inclined orbits that are thought to have been dynamically perturbed by Neptune. Further out, detached objects such as Sedna have perihelia that are far from Neptune’s influence, while some may represent the transition to the hypothesized inner Oort cloud.
@@
The remote location of these icy bodies, which shields them from the intense solar radiation that affects objects closer to the Sun, makes them valuable objects for studies of the early physical and dynamical conditions that shaped our planetary system. However, it is challenging given their relatively small size (the largest TNO is Pluto, with a radius of 1188 km (Nimmo et al. 2017)), along with their low brightness and large heliocentric distances. Moreover, the low surface temperatures of TNOs (30-50 K, e.g. Stern & Trafton (2008)) cause their thermal emission to peak in the far-infrared (FIR) region of the electromagnetic spectrum, which is blocked by Earth’s atmosphere. This makes space-based observations essential for their study. However, the stellar occultation technique enables ground-based observations and has proven to be one of the most effective methods for studying TNOs to date. This technique has significantly advanced our understanding of these distant icy objects by allowing for highly accurate size measurements, revealing the presence of atmospheres, moons, or rings. In addition, when combined with independent datasets, it helps determine their geometric albedo (e.g. Hubbard et al. (1988); Sicardy et al. (2011); Ortiz et al. (2017); Assafin et al. (2023)). Moreover, it achieves kilometric precision, making it especially useful for refining the orbital parameters through astrometry and making highly accurate shape determinations.
@@
In this work, we focus on the trans-Neptunian object (119951) 2002 KX14, discovered in 2002, which has a semi-major axis of 38.63 au, an eccentricity of 0.04, and an inclination of 0.41°. This object exhibits orbital characteristics; specifically, low eccentricity and low inclination, which are typical of the cold classical population. Members of this population are the most dynamically stable objects in the classical belt and are believed to have formed in situ, making them valuable tracers of the early formation history of the trans-Neptunian region (Delsanti & Jewitt 2006). However, the semi-major axis of 2002 KX14 places it outside the classical belt region, which is generally defined between $\sim$42 and 47 au. Some authors (e.g., Fernández-Valenzuela et al. 2021) have classified it as an inner classical object, a population that is not considered to be an extension of the cold classicals, but instead belonging to the low-inclination tail of the hot population (Vilenius et al. 2012). Unlike the cold classicals, the hot population is thought to have been dynamically excited and relocated to its current position during the migration of the giant planets. During this migration, gravitational interactions perturbed their orbits, increasing both their inclination and eccentricity. Therefore, studying 2002 KX14 is particularly valuable because its orbital properties place it near the boundary between the cold and hot classical populations, offering insights into the dynamical structure and evolutionary history of the trans-Neptunian region.
@@
Based on thermal measurements with the Herschel Space Telescope, Vilenius et al. (2012) estimated a diameter of $455\pm 27$ km and a geometric albedo ($p_{V})$ of $9.7\begin{subarray}{c}+1.4\\ -1.3\end{subarray}\%$. This diameter is close to the size where hydrostatic equilibrium may no longer be maintained (Tancredi & Favre 2008). Subsequent ground-based observations studied its photometric properties, providing values of its absolute magnitude in the R-band: $H_{R}=4.349\pm 0.124$ (Peixinho et al. 2012), $H_{R}=4.46\pm 0.01$ (Benecchi & Sheppard 2013), and in the V-band: $H_{V}=5.07\pm 0.03$ (Boehnhardt et al. 2014), $H_{V}=4.978\pm 0.017$ (Alvarez-Candal et al. 2016). Also, Alvarez-Candal et al. (2014) recorded a positive occultation, measuring a chord of $415\pm 1$ km and retrieving an equivalent diameter of $365\begin{subarray}{c}+30\\ -21\end{subarray}$ km. Subsequently, Rommel et al. (2020) obtained two additional chords of $421\pm 3$ km and $431\pm 6$ km. Despite these successful observations, due to its faint apparent magnitude above 20.5 in the V-band, with rotational variations under 0.05 mag (Benecchi & Sheppard 2013), the shape of 2002 KX14 as well as its pole and rotation period are unknown.
@@
In this study, we present the results of five positive occultations, which yield a total of 15 chords on five different dates. These observations have allowed us to determine the object’s shape and geometric albedo. Currently, only thirteen TNOs have projected shape determinations (Ortiz et al. 2012; Benedetti-Rossi et al. 2016; Ortiz et al. 2017; Dias-Oliveira et al. 2017; Benedetti-Rossi et al. 2019; Ortiz et al. 2020; Souami et al. 2020; Santos-Sanz et al. 2022; Vara-Lubiano et al. 2022; Pereira et al. 2023; Rommel et al. 2023; Kretlow et al. 2024), with 2002 KX14 being now the first [^2] Apart from Arrokoth, which was observed up close by the New Horizons mission cold classical object (if confirmed). The fact that it is close to the critical diameter beyond which hydrostatic equilibrium may no longer be maintained makes determining its shape a particularly interesting undertaking. Therefore, this study is poised to contribute to the growing body of knowledge on TNOs, laying the groundwork for future in-depth studies, providing data to enhance the statistics that allow us to contextualize this population more comprehensively, and demonstrating the effectiveness of the occultation technique in characterizing these distant objects.
@@ 2 Observations
Table 1: Occulted stars by 2002 KX14 presented in this work.
@@
 [^3] Stellar right ascension (h m s) and declination (d m s) represent the astrometric coordinates on the International Celestial Reference Frame (ICRF) of the stars, corrected of proper motion, given for the moments of the occultations.
Initial predictions for 2002 KX14 were made using the Numerical Integration of the Motion of an Asteroid (NIMA) solution (Desmars et al. 2015). Subsequently, several observational campaigns were conducted prior to each occultation using the 1.23 m telescope at the Calar Alto Observatory and the 2.0 m Liverpool Telescope located at the Roque de los Muchachos Observatory. These efforts were aimed at performing astrometry for this TNO and refining the predicted occultation shadow path, which was communicated to observers for proper deployment, in cases where mobile equipment was used. This approach enabled the observation of five positive occultations in 2020, 2022, and 2023, as described in the following sections.
2.1 Occultation A: May 26, 2020
@@
The event was observed from nine different locations, resulting in six positive detections from Borowiec (Poland), Odesa-Mayaki Observatory (Ukraine), Konkoly Observatory (Hungary), Gothard Astrophysical Observatory (Hungary), Millen-Bruch (Germany), and Nandrin (Belgium); and two negative detections: from Buelach (Switzerland) and Çukurova (Turkey). In addition, it was also observed from Tubitak Observatory (Turkey) and Marseilles (France), but the signal-to-noise ratio (S/N) of the data was too low to confirm whether the occultation occurred or not. An additional positive detection was obtained from Strasice (Czechia), but due to technical issues with the data format, it was not possible to determine the ingress and egress times with sufficient accuracy to be useful for this work. Fig. 3(a) shows the predicted occultation shadow path along with the position of the observing stations involved in this work. Table 3 expands on the information regarding locations, instrumentation, and the observers involved. Time series observations were obtained and time synchronized using Network Time Protocol (NTP) servers or global positioning systems (GPS) devices. Image acquisition started at least $\sim$5 min before and ended $\sim$5 min after the occultation time. No filters were used to maximize the S/N of the occulted star.
@@
As it is a giant star (see the Hertzsprung–Russell diagram in Fig. 2) $(a,b)=(0.669,0.223)$ according to van Belle (1999). $V$ and $K$ denote the apparent magnitudes of the star in the V-band and K-band, being $V$ = 14.750 mag and $K$ = 10.520 mag (Zacharias et al. 2004). Consequently, the angular diameter is calculated as 0.046 mas, corresponding to 1.26 km at a 2002 KX14’s geocentric distance ($\Delta$) of 38.01 au. As the central wavelength ($\lambda$) is $\sim$600 nm for our CCD/CMOS observations, using Eq. 2, the Fresnel scale of this object can be calculated as 1.3 km,
@@
2.2 Occultation B: June 23, 2022
On June 23, 2022, an occultation of the star Gaia DR3 4116601049409385472 by 2002 KX14 (Table 1), with a predicted maximum duration of 18.7 seconds, was forecasted to occur at 00:51:33 UT (hereafter, ”Occ. B”). The event was observed from four locations in Spain: the Botorrita, Javalambre, La Hita, and Calar Alto observatories. Three of these observatories, Javalambre, La Hita, and Calar Alto, reported positive detections, while Botorrita experienced technical issues (Fig. 3(b), Table 3). Time series observations were captured and synchronized using either NTP servers or GPS devices. Image acquisition started approximately $\sim$5 minutes before and concluded $\sim$5 minutes after the predicted event time. No filters were applied to optimize the S/N of the star.
@@ 2.3 Occultation C: May 17, 2023
An angular diameter of 0.046 mas of the occulted star was estimated using Eq. 1 (giant star, see Fig. 2), with $V$ = 16.47 mag and $K$ = 10.71 mag from Zacharias et al. (2004). This corresponds to 1.25 km at 2002 KX14’s geocentric distance, $\Delta$, of 37.97 au. The Fresnel scale in this case is 1.3 km. Because the minimum cycle time for these observations is 2.0 seconds, while 2002 KX14’s velocity at the time of occultation was 20.38 km/s, it translates to a distance of 40.76 km. Given that this distance is orders of magnitude greater than that derived from the star’s angular size and Fresnel diffraction effects, the stellar angular size and diffraction have a slight impact on the derivation of ingress and egress times. However, they were accounted for.
2.4 Occultation D: May 19, 2023
Two days later, on May 19, 2023, an occultation by 2002 KX14 of the star Gaia DR3 4068793394867887872 (Table 1) with a maximum duration of 21.8 seconds was predicted to happen at 03:08:49 UT (hereafter ”Occ. D”). Observations of this event were conducted from three locations in Brazil, with positive detections reported by the GOA and Sonear3 observatories; whereas in DogsHeaven, the S/N of the data was too low (Fig. 3(d), Table 3). Time series data were collected and synchronized using NTP servers. The image capture began approximately $\sim$5 min before the event and continued for about five minutes afterward. To maximize the star’s S/N, no filters were applied.
Using Eq. 1 (giant star; see Fig. 2), the occulted star’s diameter was estimated to be 0.032 mas, with magnitudes of $V$ = 16.39 mag and $K$ = 11.36 mag from Zacharias et al. (2004). At 2002 KX14’s geocentric distance of 37.96 au, this angular size corresponds to a physical diameter of 0.89 km. The Fresnel scale is again 1.3 km. Given that the minimum observation cycle time is 6.8 seconds and that 2002 KX14’s velocity during the occultation was 20.83 km/s, this timing translates to a spatial distance of 141.64 km. Thus, the ingress and egress determination uncertainty mainly comes from the photometric accuracy of the occultation light curves. Even so, the star’s angular size and Fresnel diffraction effects were included in the analysis.
2.5 Occultation E: July 01, 2023